OpenStudy (bunnykitten02):

please help Which line shows the first error in the solution? 8x-6=4x-2 (1) 8x-4x-6=4x-4x-2 (2) 4x-6=2 (3) 4x-6+6=2+6 (4) 4x=8 (5) 4x/4= 8/4 (6) x=2 (7) .A. line 3 B. line 4 C. line 5 D. line 6

OpenStudy (michele_laino):

hint: at the second step, I can write this: \(4x-4x-2=-2\)

OpenStudy (michele_laino):

it is the right side

OpenStudy (bunnykitten02):

well my dad said it was line 2 but it isnt an answer choice

OpenStudy (michele_laino):

your dad is right! since \(4x-4x-2=-2\) and not \(2\), so the second line is: \(4x-6=-2\)

OpenStudy (bunnykitten02):

but it isnt an answer choice so what do i do

OpenStudy (michele_laino):

maybe there is a typo in the numeration is wrong, namely the first line is: \(8x-6=4x-2\), so the second line is \(8x-4x-6=4x-4x-2\) and the third line is \(4x-6=2 \), and so on...

OpenStudy (michele_laino):

oops.. I meant maybe there is a typo in the numeration

OpenStudy (michele_laino):

in other words, here is the right numeration: (1) 8x-6=4x-2 (2) 8x-4x-6=4x-4x-2 (3) 4x-6=2 (4) 4x-6+6=2+6 (5) 4x=8 (6) 4x/4= 8/4 (7) x=2 (7)

OpenStudy (bunnykitten02):

k so whats the answer

OpenStudy (michele_laino):

hint: we see that there is an error on line (3)

OpenStudy (bunnykitten02):

there is?

OpenStudy (michele_laino):

it is option A, am I right?

OpenStudy (bunnykitten02):

i beleve so and my dad agrees thank you

OpenStudy (michele_laino):

:)